OpenStudy (anonymous):

integration sin^2 theta / cos^2 theta * sin^2 theta

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ \sin^2 \theta }{ \cos^2 \theta \sin^2 \theta }\]

OpenStudy (anonymous):

can i cancel out the all the squre?

OpenStudy (amoodarya):

first simplify sin2(th) then note that 1/cos^2(th) = 1+tan^2(th) so you have int(1+yan^2(th))d(th) now its easy

OpenStudy (amoodarya):

answer is tan(theta) +c(constant)

OpenStudy (anonymous):

after the simplify, i got 1/cos^2(th) and i changed this as int sec^2(th) and now, its tan(th) +c ?

OpenStudy (anonymous):

you said, 1/cos^2(th) = 1+tan^2(th) --------> this is also equal to sec^2(th) right?

OpenStudy (anonymous):

i got the same answer as you

OpenStudy (amoodarya):

yes 1/cos^2(th) = 1+tan^2(th)=sec^2(th)

OpenStudy (anonymous):

but then i have limit 0 to 2pi ---> so tan(th) is 0?

OpenStudy (anonymous):

* int than(th) is 0?

OpenStudy (amoodarya):

wait a mooment

OpenStudy (amoodarya):

no the answer is not 0 because function is positive in(0-2pi)

OpenStudy (amoodarya):

answer of this int (0 - 2pi) is +infiity

OpenStudy (anonymous):

hmm its hard.. hahaha

OpenStudy (anonymous):

actually this problem is from double integrals in polar coordinates and that i asked the first was just the part of it

OpenStudy (anonymous):

is that mean limit 0 to 2pi should be changed?

OpenStudy (amoodarya):

is it your home work ? area goes to infinity

OpenStudy (amoodarya):

because of the vertical retricempote x=pi/2, 3pi/2